For what value of b in the equation 2x + 3b – 3 = 4x – 3b will the solution for x be 3?

Answer by sweet_vry_sweet_funky(1)

2x + 3b - 3 = 4x - 3b
or,3b + 3b = 4x - 2x + 3
or,6b = 2x + 3.
(now, considering x=3),
or,6b = 2*3 + 3
or,6b = 9
or,b = 9/6
or,b = 3/2
ANSWER:3/2